# Onboard Satellite Image Classification for Earth Observation: A Comparative Study of ViT Models

### v2

@@ Abstract
Remote sensing (RS) image classification (IC) is a critical component of Earth observation (EO) systems, traditionally dominated by convolutional neural networks (CNNs) and other deep learning (DL) techniques. However, the advent of Transformer-based architectures and large-scale pre-trained models has significantly shifted the trend by offering enhanced performance and efficiency. Hence, this study focuses on identifying the most effective pre-trained model for land use classification in onboard satellite processing, emphasizing achieving high accuracy, computational efficiency, and robustness against noisy data—conditions commonly encountered during satellite-based inference. Through extensive experimentation, we compare the performance of traditional CNN-based, ResNet-based, and various pre-trained vision Transformer models. Our findings demonstrate that pre-trained Vision Transformer (ViT) models, particularly MobileViTV2 and EfficientViT-M2, outperform models trained from scratch in terms of accuracy and efficiency. These models achieve high performance with reduced computational requirements and exhibit greater resilience during inference under noisy conditions. While MobileViTV2 has excelled on clean validation data, EfficientViT-M2 has proved to be more robust when handling noise, making it the most suitable model for onboard satellite EO tasks. Our experimental results demonstrate that EfficientViT-M2 is the optimal choice for reliable and efficient RS-IC in satellite operations, achieving 98.76% of accuracy, precision, and recall. Precisely, EfficientViT-M2 delivers the highest performance across all metrics, excels in training efficiency (1,000s) and inference time (10s), and demonstrates greater robustness (overall robustness score of 0.79). Consequently, EfficientViT-M2 consumes 63.93% less power than MobileViTV2 (79.23 W) and 73.26% less power than SwinTransformer (108.90 W). This highlights its significant advantage in energy efficiency.
@@ I Introduction
However, recent advances have shifted towards Transformer-based architectures, surpassing conventional CNNs in several IC tasks, including those in RS applications [5, 6]. The rise of large-scale pre-trained models has marked a new era in artificial intelligence (AI), achieving breakthroughs across various domains, particularly in natural language processing with models like BERT [7]. This paradigm shift extends into computer vision, where pre-trained ViTsize models demonstrate superior performance across a wide range of tasks [8, 9]. These models benefit from extensive pre-training on large datasets, allowing them to learn rich contextual representations that can be fine-tuned for specific tasks with minimal additional training. The success of these models is primarily attributed to advancements in computational power, the availability of vast amounts of data, and innovations in model architecture and efficiency [10]. While pre-trained ViT models hold great potential for enhancing classification performance and reducing computational demands during inference, their application in onboard satellite systems (OSS) is constrained due to such large-scale models’ high power consumption and computing resource requirements. Implementing these models for OSS real-time decision-making remains a significant challenge.
Furthermore, a major operational bottleneck for LEO satellites is their dependence on ground stations for data transmission, leading to limited communication opportunities. The short visibility window results in prolonged connectivity loss, delaying access to critical data and hindering timely responses [11].
@@
Therefore, most onboard neural networks (NNs) are designed to be lightweight, prioritizing computational efficiency over model complexity. For example, the $\Phi$-Sat-1 mission employed the CloudScout model, a CNN-based NN, for onboard image segmentation on the Intel Movidius Myriad 2 vision processing unit (VPU), making it the first instance of DL deployed on a satellite [14]. Similarly, the $\Phi$-Sat-2 mission implemented a convolutional autoencoder for image compression to minimize data transmission needs [15] and tested on three different hardware, including graphic processing unit (GPU) NVIDIA GeForce GTX 1650 VPU Myriad 2, and central processing unit (CPU) Intel Core i7-6700. These examples highlight the practical limitations in deploying more complex models, such as Transformer-based architectures and onboard satellites, due to restricted power budgets and limited processing resources. Consequently, while CNN-based models can achieve reasonable performance, they cannot often capture the intricate patterns and contextual dependencies required for more advanced tasks, limiting their effectiveness in real-world remote sensing applications. One promising approach to onboard processing is using neuromorphic processors, which rely on converting traditional NNs into Spiking NNs (SNNs). Currently, this conversion is best supported for CNN-based models, as demonstrated in recent studies on energy-efficient neuromorphic computing for space applications [16, 17]. While these lightweight CNN-based models are advantageous due to their low power consumption and reduced computational requirements, they often fall short in terms of prediction accuracy and robustness compared to more complex Transformer-based models, which are better suited for capturing intricate patterns and contextual dependencies in RS data.
@@
  • Adaptation of pre-trained models for resource-constrained environments: We demonstrate how large-scale, pre-trained Transformer models can be effectively adapted for onboard satellite processing, providing a new framework for deploying complex AI models in power- and computation-limited environments while maintaining high accuracy and efficiency.
  • Optimizing real-time decision-making in satellite operations: EfficientViT-M2 is identified as the optimal model for onboard inference, demonstrating robustness with noisy sensor data and enabling autonomous, real-time decisions. In scenarios like forest fires or oil spills, the satellite can autonomously repoint its sensors or switch to infrared imaging for better visibility without waiting for ground-based instructions, significantly improving disaster response and environmental monitoring efficiency.
  • Bridging the gap between ground and onboard processing: By conducting rigorous ground-based evaluations of state-of-the-art models and then translating these findings to onboard implementation, we bridge the gap between theoretical AI research and practical applications in space, highlighting a pathway for deploying advanced AI solutions to enhance operational efficiency and autonomy in satellite EO missions.
The remainder of this paper is organized as follows. Section II provides a brief survey of related works, serving as the basis for highlighting this paper’s contributions. Section III presents the methodologies, including the dataset, machine learning (ML) models, and evaluation metrics employed in this comprehensive study. The experimental setup is explained in Section IV, followed by the numerical results and discussion in Section V. Finally, Section VI concludes the paper. Reproducible codes for data augmentation, noisy data, training, and inference are available from our shared Github repository 111 https://github.com/ltdung/SnT-SENTRY.
@@ III-A Dataset
The public benchmark for land use and land cover classification dataset, EuroSAT [32], will be used in this study. The EuroSAT dataset is a large-scale benchmark dataset designed explicitly for land use and land cover classification derived from Sentinel-2 satellite imagery. It comprises 27,000 geo-referenced labeled images, each measuring 64x64 pixels and spanning 13 spectral bands. The dataset is categorized into 10 classes. Each class contains 2000–3000 images, including industrial buildings, residential buildings, annual crops, permanent crops, rivers, seas and lakes, Herbaceous vegetation, highways, pastures, and forests across Europe. Due to its compact image size and diverse class representation, EuroSAT is particularly well-suited for developing and evaluating DL models intended for onboard satellite processing in EO missions. This makes it a valuable resource for applications such as real-time environmental monitoring, disaster response, and precision agriculture, where in-situ processing capabilities are crucial for timely decision-making.
@@ V Results and Discussions
TABLE III: Experimental results for the best model’s performance
@@
From the results presented in Tables II, and III, and Fig. 3, and 4 we confirm that MobileViT and EfficientViT are the two best models for IC. These models excel due to their lower computational complexity, shorter training and inference times, and superior performance across all evaluation metrics. Notably, compared to state-of-the-art models, as shown in Table IV, the exceptional performance of EfficientViT-M2 and MobileViTV2 becomes evident. In the EuroSAT IC task, MobileViTV2 achieves the highest scores across all evaluation metrics, outperforming all other models in the comparison. EfficientViT-M2 also performs exceptionally well, securing the second-highest results. These findings highlight the effectiveness of both EfficientViT-M2 and MobileViTV2 in achieving top-tier performance with relatively lower computational complexity and faster training and inference times, as observed in our experiment results. This confirms that these two models are currently the best choices for high-performance IC tasks, providing a significant advantage over existing methods.
Additionally, Fig. 5 shows the confusion matrices for EfficientViT-M2 and MobileViTV2, highlighting their exceptional performance in IC. Both models exhibit vital accuracy across all classes, with most predictions aligning closely with the true labels. In the confusion matrix for EfficientViT-M2, we observe that classes like AnnualCrop, Forest, PermanentCrop, and SeaLake are classified with nearly perfect accuracy, showing minimal misclassification. Similarly, the confusion matrix for MobileViTV2 shows even fewer errors, with several classes like Forest, PermanentCrop, and SeaLake being classified almost flawlessly. A few minor misclassifications are observed in both models, particularly in categories like Pasture and River, but these are relatively small compared to the overall high performance. These matrices underscore the effectiveness of both models in handling complex IC tasks, with MobileViTV2 slightly outperforming EfficientViT-M2 in terms of fewer misclassifications, further affirming its position as the top-performing model.
Figure 5: Confusion matrix from MobileViTV2 (top) and EfficientViT-M2 (bottom) performance.
TABLE IV: Performance comparison with the-state-of-the-art models (Best run)
@@
TABLE V: MobileViTV2 Performance with Different Augmentation Techniques
@@
TABLE VI: EfficientViT Performance with Different Augmentation Techniques
@@
The experiment results, as presented in Tables V and VI, reveal that applying data augmentation techniques during training and inference generally leads to a slight reduction in model performance when evaluated on clean test data. MobileViTV2’s and EfficientViT-M2’s baseline performance—achieved without augmentation—yield the highest accuracy. However, among the various augmentation techniques tested, RandErasing for both models demonstrates the most effective results, providing performance metrics (98.73% for MobileViTV2 and 98.22% for EfficientViT-M2) that are closest to the baseline. This indicates that RandErasing introduces sufficient variability to enhance robustness without significantly compromising the models’ ability to perform well on clean data. Although there is a slight decrease compared to the baseline, using RandErasing offers a good balance, maintaining high accuracy while preparing the models to handle real-world scenarios with noisy inputs.
@@
Furthermore, the feasibility of deploying the EfficientViT-M2 model in practical applications, mainly OSS, is underscored by its relatively low computational complexity and compact size. With an estimated total size of 38.19 MB and a computational complexity of 203.53 MFLOPs, EfficientViT-M2 offers a significant advantage over earlier models used in pioneering satellite missions. For instance, the $\Phi$-Sat-1 mission [14], which utilized the CloudScout model for image segmentation, required approximately 4.67 GFLOPs and had an estimated total size of 123.98 MB. In contrast, EfficientViT-M2 requires substantially less computational power and storage, making it a more efficient choice for similar tasks. Additionally, compared to the $\Phi$-Sat-2 mission [15], which employed a convolutional autoencoder with 81.2 MFLOPs for vision processing on an Intel Movidius Myriad 2 VPU, EfficientViT-M2’s slightly higher complexity is justified by its enhanced robustness and performance capabilities. While the $\Phi$-Sat-2 model took nearly 11 seconds to compress an image with a compression ratio of 8 on a VPU, EfficientViT-M2’s processing requirements are more aligned with modern GPU capabilities, completing complex tasks much faster and with greater accuracy. These factors make EfficientViT-M2 a powerful model for IC and a practical and feasible option for real-world deployment in resource-constrained environments like satellite missions.
@@ VI Conclusions
Finally, the study is limited by its focus on single-modality data, potentially restricting its ability to leverage the expanding wealth of multimodal EO data. Future research should explore the potential of transformer-based approaches within a multimodal DL framework to utilize better and integrate diverse types of EO data [53, 54].

### v3

@@ Abstract
Remote sensing (RS) image classification (IC) is a critical component of Earth observation (EO) systems, traditionally dominated by convolutional neural networks (CNNs) and other deep learning (DL) techniques. However, the advent of Transformer-based architectures and large-scale pre-trained models has significantly shifted the trend by offering enhanced performance and efficiency. Hence, this study focuses on identifying the most effective pre-trained model for land use classification in onboard satellite processing, emphasizing achieving high accuracy, computational efficiency, and robustness against noisy data—conditions commonly encountered during satellite-based inference. Through extensive experimentation, we compare the performance of traditional CNN-based, ResNet-based, and various pre-trained vision Transformer models. Our findings demonstrate that pre-trained Vision Transformer (ViT) models, particularly MobileViTV2 and EfficientViT-M2, outperform models trained from scratch in terms of accuracy and efficiency. These models achieve high performance with reduced computational requirements and exhibit greater resilience during inference under noisy conditions. While MobileViTV2 has excelled on clean validation data, EfficientViT-M2 has proved to be more robust when handling noise, making it the most suitable model for onboard satellite EO tasks. Our experimental results demonstrate that EfficientViT-M2 is the optimal choice for reliable and efficient RS-IC in satellite operations, achieving 98.76% of accuracy, precision, and recall. Precisely, EfficientViT-M2 delivers the highest performance across all metrics, excels in training efficiency (1,000s) and inference time (10s), and demonstrates greater robustness (overall robustness score of 0.79). Consequently, EfficientViT-M2 consumes 63.93% less power than MobileViTV2 (79.23 W) and 73.26% less power than SwinTransformer (108.90 W). This highlights its significant advantage in energy efficiency. Reproducible codes for data augmentation, noisy data, training, and inference are available from our shared Github repository 111 https://github.com/ltdung/SnT-SENTRY.
@@ I Introduction
However, recent advances have shifted towards Transformer-based architectures, surpassing conventional CNNs in several IC tasks, including those in RS applications [5, 6]. The rise of large-scale pre-trained models has marked a new era in artificial intelligence (AI), achieving breakthroughs across various domains, particularly in natural language processing with models like BERT [7]. This paradigm shift extends into computer vision, where pre-trained ViT models demonstrate superior performance across a wide range of tasks [8, 9]. These models benefit from extensive pre-training on large datasets, allowing them to learn rich contextual representations that can be fine-tuned for specific tasks with minimal additional training. The success of these models is primarily attributed to advancements in computational power, the availability of vast amounts of data, and innovations in model architecture and efficiency [10]. While pre-trained ViT models hold great potential for enhancing classification performance and reducing computational demands during inference, their application in onboard satellite systems (OSS) is constrained due to such large-scale models’ high power consumption and computing resource requirements. Implementing these models for OSS real-time decision-making remains challenging. Furthermore, a major operational bottleneck for LEO satellites is their dependence on ground stations for data transmission, leading to limited communication opportunities. The short visibility window results in prolonged connectivity loss, delaying access to critical data and hindering timely responses [11].
@@
Therefore, most onboard neural networks (NNs) are designed to be lightweight, prioritizing computational efficiency over model complexity. For example, the $\Phi$-Sat-1 mission employed the CloudScout model, a CNN-based NN, for onboard image segmentation on the Intel Movidius Myriad 2 vision processing unit (VPU), making it the first instance of DL deployed on a satellite [14]. Similarly, the $\Phi$-Sat-2 mission implemented a convolutional autoencoder for image compression to minimize data transmission needs [15] and tested on three different hardware, including graphic processing unit (GPU) NVIDIA GeForce GTX 1650, VPU Myriad 2, and central processing unit (CPU) Intel Core i7-6700. These examples highlight the practical limitations in deploying more complex models, such as Transformer-based architectures and onboard satellites, due to restricted power budgets and limited processing resources. Consequently, while CNN-based models can achieve reasonable performance, they cannot often capture the intricate patterns and contextual dependencies required for more advanced tasks, limiting their effectiveness in RS.
The computational demands of ViTs - especially in model complexity and resource consumption - are one of the primary challenges in their onboard deployment. Unlike CNNs, which are relatively lightweight and optimized for image processing, ViTs require substantial computational resources, particularly for large models with numerous parameters and attention mechanisms. A recent study proposed a solution to this challenge by introducing RepSViT, a lightweight ViT-based model designed for onboard image processing. With only 3.77 million parameters and a computational cost of just 0.6 GFLOPs, RepSViT demonstrates that it is possible to deploy a compact version of ViT for satellite missions, while still capturing some of the advantages of Transformer architectures. However, such models, though promising, still face challenges when it comes to processing power and energy efficiency, making them less suited for complex tasks unless further model compression and optimization techniques are applied [16]. One promising approach to onboard processing is using neuromorphic processors, which rely on converting traditional NNs into Spiking NNs (SNNs). Currently, this conversion is best supported for CNN-based models, as demonstrated in recent studies on energy-efficient neuromorphic computing for space applications [16, 17, 18]. While these lightweight CNN-based models are advantageous due to their low power consumption and reduced computational requirements, they often fall short in prediction accuracy and robustness compared to more complex Transformer-based models, better suited for capturing patterns and contextual dependencies in RS data.
@@
  • Energy‑aware pretrained ViT selection: A systematic exploration of pre‑trained ViTs highlights EfficientViT‑M2 as the best accuracy-per-power consumption option for onboard EO, satisfying the constrained power budgets of onboard processing.
  • Comprehensive robustness validation: Inference reliability is proven under unseen perturbations, across multiple benchmark EO datasets, and within a SatCom‑in‑the‑loop framework that injects DVB‑S2(X) channel impairments - demonstrating stable performance from sensor capture to downlink.
The remainder of this paper is organized as follows. Section II provides a brief survey of related works, serving as the basis for highlighting this paper’s contributions. Section III presents the methodologies, including the dataset, machine learning (ML) models, and evaluation metrics employed in this comprehensive study. The experimental setup is explained in Section IV, followed by the numerical results and discussion in Section V. Finally, Section VI concludes the paper.
@@ III-A Dataset
The EuroSAT dataset [33], a public benchmark for land use and land cover classification, will be used in this study. Specifically designed for land use and land cover classification, the EuroSAT dataset is derived from Sentinel-2 satellite imagery and comprises 27,000 geo-referenced labeled images, each measuring 64x64 pixels and spanning 13 spectral bands. It is categorized into 10 classes, including industrial buildings, residential buildings, annual crops, permanent crops, rivers, seas and lakes, herbaceous vegetation, highways, pastures, and forests across Europe. With its compact image size and diverse class representation, EuroSAT is particularly well-suited for developing and evaluating deep learning models intended for onboard satellite processing in EO missions. In addition, the models will also be tested on the PatternNet dataset [34], which focuses on image retrieval using 30,400 high-resolution, 256x256 pixel images spanning 38 categories. The dataset’s broad and diverse coverage enables the development of retrieval-focused approaches for addressing more complex, fine-grained remote sensing challenges. It is particularly valuable for applications such as real-time environmental monitoring and precision agriculture, where onboard processing capabilities are essential for timely decision-making.
@@ V Results and Discussions
TABLE III: Experimental results for the best model’s performance on EuroSat
@@
From the results presented in Tables II, and III, and Fig. 3, and 4 we confirm that MobileViT and EfficientViT are the two best models for IC on EuroSat. These models excel due to their lower computational complexity, shorter training and inference times, and superior performance across all evaluation metrics. Notably, compared to state-of-the-art models, as shown in Table IV, the exceptional performance of EfficientViT-M2 and MobileViTV2 becomes evident. In the EuroSAT IC task, MobileViTV2 achieves the highest scores across all evaluation metrics, outperforming all other models in the comparison. EfficientViT-M2 also performs exceptionally well, securing the second-highest results. These findings highlight the effectiveness of both EfficientViT-M2 and MobileViTV2 in achieving top-tier performance with relatively lower computational complexity and faster training and inference times, as observed in our experiment results. This confirms that these two models are currently the best choices for high-performance IC tasks, providing a significant advantage over existing methods.
Additionally, Fig. 5 shows the confusion matrices for EfficientViT-M2 and MobileViTV2, highlighting their exceptional performance in IC on EuroSat. Both models exhibit vital accuracy across all classes, with most predictions aligning closely with the true labels. In the confusion matrix for EfficientViT-M2, we observe that classes like AnnualCrop, Forest, PermanentCrop, and SeaLake are classified with nearly perfect accuracy, showing minimal misclassification. Similarly, the confusion matrix for MobileViTV2 shows even fewer errors, with several classes like Forest, PermanentCrop, and SeaLake being classified almost flawlessly. A few minor misclassifications are observed in both models, particularly in categories like Pasture and River, but these are relatively small compared to the overall high performance. These matrices underscore the effectiveness of both models in handling complex IC tasks, with MobileViTV2 slightly outperforming EfficientViT-M2 in terms of fewer misclassifications, further affirming its position as the top-performing model.
TABLE IV: Performance comparison with the-state-of-the-art models (Top-1) on EuroSat.
@@
Furthermore, Table V and the accompanying confusion matrix, Fig. 6, clearly demonstrate that EfficientViT‑M2 achieves near‑perfect scene recognition on PatternNet. With 99.52% Top‑1 accuracy, precision, and recall, the model sits within 0.14 percentage points of the best‑reported score while outperforming all remaining baselines, including Swin‑Transformer (99.23%) and ResNet‑50 (98.23%) [54]. The confusion matrix perfectly classifies 38 classes, confirming exceptionally consistent predictions across every land‑use category. This combination of accuracy, uniform class‑level reliability, and the architecture’s lightweight design, again highlights EfficientViT‑M2 as an excellent choice for onboard EO applications where both performance and computational efficiency are critical.
TABLE V: Performance comparison with the-state-of-the-art models (Top-1) on PatternNet.
Models Accuracy Precision Recall
EfficientViT-M2 99.52 99.52 99.52
MobileViTV2 99.66 99.66 99.66
EfficientViT-b2 99.54 99.54 99.53
ResNet-DINO 96.01 96.01 96.00
SwinTransformer 99.23 99.23 99.23
ResNet-50 [54] 98.23 97.95 x
          Bold denotes the best values.
          Italic and underline denote the second best values.
          Bold and underline denote the third best values.
Figure 5: Confusion matrix from MobileViTV2 (top) and EfficientViT-M2 (bottom) performance on EuroSat.
Figure 6: Confusion matrix of EfficientViT-M2 on PatternNet.
TABLE VI: MobileViTV2 Performance with Different Augmentation Techniques
@@
TABLE VII: EfficientViT Performance with Different Augmentation Techniques
@@
The experiment results, as presented in Tables VI and VII, reveal that applying data augmentation techniques during training and inference generally leads to a slight reduction in model performance when evaluated on clean test data. MobileViTV2’s and EfficientViT-M2’s baseline performance—achieved without augmentation—yield the highest accuracy. However, among the various augmentation techniques tested, RandErasing for both models demonstrates the most effective results, providing performance metrics (98.73% for MobileViTV2 and 98.22% for EfficientViT-M2) that are closest to the baseline. This indicates that RandErasing introduces sufficient variability to enhance robustness without significantly compromising the models’ ability to perform well on clean data. Although there is a slight decrease compared to the baseline, using RandErasing offers a good balance, maintaining high accuracy while preparing the models to handle real-world scenarios with noisy inputs.
@@
Furthermore, the feasibility of deploying the EfficientViT-M2 model in practical applications, mainly OSS, is underscored by its relatively low computational complexity and compact size. With an estimated total size of 38.19 MB and a computational complexity of 203.53 MFLOPs, EfficientViT-M2 offers a significant advantage over earlier models used in pioneering satellite missions. For instance, the $\Phi$-Sat-1 mission [14], which utilized the CloudScout model for image segmentation, required approximately 4.67 GFLOPs and had an estimated total size of 123.98 MB. In contrast, EfficientViT-M2 requires substantially less computational power and storage, making it a more efficient choice for similar tasks. Additionally, compared to the $\Phi$-Sat-2 mission [15], which employed a convolutional autoencoder with 81.2 MFLOPs for vision processing on an Intel Movidius Myriad 2 VPU, EfficientViT-M2’s slightly higher complexity is justified by its enhanced robustness and performance capabilities. While the $\Phi$-Sat-2 model took nearly 11 seconds to compress an image with a compression ratio of 8 on a VPU, EfficientViT-M2’s processing requirements are more aligned with modern GPU capabilities, completing complex tasks much faster and with greater accuracy. Furthermore, the RepSViT model [16] utilizes 3.77 million parameters and consumes 600 MFLOPs of computational costs for onboard image processing, but EfficientViT-M2 offers a $\approx$3x reduction in complexity compared to RepSViT. This lower computational cost, coupled with its efficient architecture, positions EfficientViT-M2 as a promising model for satellite EO, offering the necessary balance between performance and resource efficiency for real-time processing in space environments. These factors make EfficientViT-M2 a powerful model for IC and a practical and feasible option for real-world deployment in resource-constrained environments like satellite missions.
Lastly, the robustness analysis was extended beyond synthetic perturbations by subjecting EfficientViT-M2 to an end‑to‑end satellite communications (SatCom) workflow that emulates the DVB‑S2(X) downlink with realistic channel coding, RF impairments, and compression losses. EfficientViT demonstrated highly efficient and robust performance for EO-IC under realistic SatCom conditions, explicitly using the DVB-S2(X) transmission standard, incorporating typical channel impairments. By systematically evaluating EfficientViT across varying image compression levels and practical channel SNR-to-Shannon-capacity ratios, the model consistently achieved high classification accuracy, remaining above 97% under optimal transmission scenarios and degrading gradually as source and channel loss increased. The consistently high accuracy from empirical results across the full SatCom operating, modeled by an exponential curve fitting framework, reveals that EfficientViT-M2 maintains strong task-oriented effectiveness and reliability for EO applications in bandwidth and noise-constrained SatCom environments, validating its practicality for real-world deployments [55].
@@ VI Conclusions
Finally, the study is limited by its focus on single-modality data, potentially restricting its ability to leverage the expanding wealth of multimodal EO data. Future research should explore the Transformer-based approaches within a multimodal DL framework to utilize better and integrate diverse types of EO data [57, 58]. Diffusion models also warrant investigation, as their iterative denoising process can generate high‑fidelity, cross‑modal representations that boost downstream performance [59]. Moreover, lightweight state‑space sequence models offer a lower‑complexity alternative to standard Transformers while still capturing long‑range dependencies [60], making them well‑suited to resource‑constrained EO deployments.